Steel train rails on LRT2 are laid in 10.0-m-long segments placed end to end. The rails are laid on a day when their temperature is 28.0°C. (a) How much space must be left between adjacent rails if they are just to touch on a summer day when their temperature is 39.0°C? (b) If the rails are





originally laid in contact, what is the stress in them on a summer day when their temperature is





39.0°C?

1
Expert's answer
2022-01-11T06:03:39-0500

Given:

"l_0=10\rm \: m"

"t_0=28.0^\circ \rm C"

"t=39.0^\circ \rm C"

"\alpha= 11.5*10^{−6}\:\rm1/°C"

"E=200\:\rm GPa"


(a) the linear exspansion equation says

"l=l_0(1+\alpha (t-t_0))"

So

"\Delta l=l_0\alpha (t-t_0)\\
=10*11.5*10^{−6}(39.0-28.0)=\rm 0.0077\: m"

(b) the stress is given by

"\sigma=\epsilon E=\frac{\Delta l}{l_0}E\\
=\frac{0.0077}{10}*200*10^9=1.54*10^8\:\rm Pa"